why does negative 1 as an exponent give you 1/5 or .5?
Answer by stanbon(48569)

-1 is an instruction.
It means "invert".
So 5^-1 = 1/5
-----------------
(2/3)^-1 = 3/2
-------------------
Cheers,
Stan H.